one machine can complete a job in 10minutes. if the same jobis done byh thismachine and an older machine working together, the job can be completed in 6 minutes. how long would it take the older machine to do the job alone?

Answer by mananth(16946)

one machine can complete a job in 10minutes. if the same jobis done byh thismachine and an older machine working together, the job can be completed in 6 minutes. how long would it take the older machine to do the job alone?
...
Machine A 6 minutes
1/6 of job in1 minute
..
Machine B + machine A = 6 minutes
they do 1/6 of the job in 1 minute
..
Job done by B in 1 minute =
1/6 -1/10
=(10-6)/60
=4/60
=1/15
...
so Machine B will take 15 minutes to do the same job